OpenStudy (anonymous):

Please help! Change 2(cos210 degrees+isin210 degrees) to rectangular (algebraic) form. a. -1+isqrt3 b. -1-isqrt3 c. -sqrt3+i d. -sqrt3-i

jimthompson5910 (jim_thompson5910):

Hint: x = r*cos(theta) y = r*sin(theta)

jimthompson5910 (jim_thompson5910):

if you let the y axis be the imaginary axis, then the number x+y*i represents the point (x,y) in the cartesian plane

OpenStudy (anonymous):

:/ Can you help me solve it step by step?

jimthompson5910 (jim_thompson5910):

Let me do a similar example and hopefully that will help

jimthompson5910 (jim_thompson5910):

Example Change 3(cos60 degrees+isin60 degrees) to rectangular (algebraic) form.

jimthompson5910 (jim_thompson5910):

In this example, r = 3 and theta = 60

jimthompson5910 (jim_thompson5910):

So x = r*cos(theta) x = 3*cos(60) x = 3*(1/2) x = 3/2 --------------- and y = r*sin(theta) y = 3*sin(60) y = 3*(sqrt(3)/2) y = (3/2)*sqrt(3) Which means 3(cos60 degrees+isin60 degrees) converts to 3/2 + (3/2)*sqrt(3)*i